# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 1.4. Structure of the article
In Section 2 we gather all the preliminary facts that we need about the structure of the group $\mathrm{G}$, its representation theory, its parabolic subgroups, and about its associated locally symmetric spaces, i.e. the Picard surfaces $\mathrm{S}_{K}$, together with the variations of Hodge structures carried by them.
In Section 3, we set up our notations for algebraic Hecke characters. Moreover, we recall the basic facts about their $L$-functions and their associated motives that we need in order to formulate the Bloch-Beilinson conjecture and to explain how Conjecture 1 follows from them.
In Section 4, we recall the structure of the Baily-Borel and Borel-Serre compactifications of Picard modular surfaces, Moreover, we revisit in the detail the computation of boundary cohomology of Picard surfaces and of its associated Hodge structure, using $\mathrm{S}^{*}_{K}$ ([1]), adding the complements we need. We proceed then by recalling in detail the contents of Harder’s paper [17] on which we rely crucially: in subsection 4.4 we record the description of boundary cohomology, using $\bar{\mathrm{S}}_{K}$, in terms of induced representations from Hecke characters, and in Section 5 we explain the description of Eisenstein cohomology, where the $L$-functions of the aforementioned Hecke characters play an essential role. We conclude the section with the construction of the first piece of our extensions, the space $\partial H^{2}(\phi)$.
In Section 6, we explain how to use Rogawski’s work in order to construct the second piece needed for our extensions, i.e. the space $H^{2}(\phi)_{!}$. The sought-for extensions are then constructed in Section 7, whose final subsection is devoted to showing how our extensions fit into suitable biextensions. These, as explained before, will provide the crucial input for our future work towards the proof of Conjecture 1.
@@ 2. Preliminaries
The first part of this section fixes notation about the ground field $F$ and the algebraic group $\mathrm{G}$ that will be in force throughout the whole paper. In the following parts, we discuss the root system and Weyl group of $\mathrm{G}$, we fix a parametrization of its irreducible representations, and we explain the structure of its parabolic subgroups. Finally, we introduce the main geometric objects of interest: the locally symmetric spaces attached to $\mathrm{G}$, their structure of Shimura varieties, and the variation of Hodge structure associated to representations of $\mathrm{G}$.
@@ 2.1. Basic notations and conventions
The symbols ${\mathbb{A}}_{{\mathbb{K}}}$ and ${\mathbb{I}}_{{\mathbb{K}}}$ will denote respectively adèles and idèles over a number field ${\mathbb{K}}$ (no subscript if ${\mathbb{K}}={\mathbb{Q}}$). The quadratic character of ${\mathbb{I}}_{F}$ associated by class field theory to the extension $F|{\mathbb{Q}}$ will be denoted by $\omega_{F|{\mathbb{Q}}}$.
@@
There always exists a $F$-basis of $V$ with respect to which $J$ is represented by a matrix of the form
@@
$\left(\begin{array}[]{ccc}1&&\\ &1&\\ &&-1\end{array}\right)$
@@
(3) $\gamma:=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}&&\frac{1}{\sqrt{2}}\\ &\sqrt{\beta}&\\ \frac{1}{\sqrt{2}}&&-\frac{1}{\sqrt{2}}\end{array}\right)$
@@
$\mathrm{G}(R)=\{g\in\mathop{\rm GL}\nolimits_{F\otimes_{{\mathbb{Q}}}R}(V\otimes_{{\mathbb{Q}}}R)|\ \exists\ \nu(g)\in R^{\times},J(g\cdot,g\cdot)=\nu(g)J(\cdot,\cdot)\}$
@@
$\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}:=\mathop{\rm SU}\nolimits(V,J)$
@@ 2.2. Structure Theory of $\mathrm{G}$
In order to describe the structure of $\mathrm{G}$, recall that $V_{F}$ decomposes as
@@
$V^{+}:=\{v\in V\otimes_{{\mathbb{Q}}}F\ |\ (x\otimes 1)v=(1\otimes x)v\ \forall x\in F\}\,$
@@
(8) $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}\simeq\mathop{\rm SL}\nolimits(V_{F})$
@@
and call ${\mathcal{B}}$ and $\mathcal{T}$ the standard (upper triangular) Borel and maximal torus of $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathop{\rm GL}\nolimits_{3,F}$. Then,
@@
is a maximal torus, that we call standard. It is isogenous to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}\times\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$. Under the isomorphism
@@
(9) $\mathrm{T}_{F}(F)\simeq\left\{(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f)|a,b,q,f\in F^{\times}\right\}$
@@
$\mathrm{T}^{\prime}:=\mathrm{T}\cap\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$
It is isomorphic to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$, and we fix the isomorphism to be the one described on ${\mathbb{Q}}$-points by [^2] This is the same convention as in [17, p. 573].
(10) $\mathrm{T}^{\prime}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)|a\in F^{\times}\right\}$
@@
(11) $\mathrm{T}^{s}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}t&&\\ &1&\\ &&t^{-1}\end{array}\right)\ |\ t\in{\mathbb{Q}}^{\times}\right\}$
@@ 2.3. Root System
$k_{1}\geq k_{2}\geq 0,\ c=k_{1}+k_{2}\pmod{2},\ r=\frac{c+k_{1}+k_{2}}{2}\pmod{2}$
@@
(12) $\left(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f\right)\mapsto a^{k_{1}-k_{2}}b^{k_{2}}q^{\frac{c-(k_{1}+k_{2})}{2}}\cdot f^{-\frac{1}{2}(r+\frac{3c-(k_{1}+k_{2})}{2})}$
@@
and the group of characters of the maximal torus $\mathrm{T}_{F}^{\prime}$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$. The isomorphism is provided by the natural factorization of characters of the above shape through $\mathrm{T}^{\prime}_{F}$. We still call the resulting character $\lambda$ and we denote it by $(k_{1},k_{2})$.
The root system $\Phi$ of the split, semisimple group $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}$ is of type $A_{2}$. Under the previous identifications, the system $\Delta$ of simple roots is the set $\{(1,-1),(1,2)\}$ and the set $\Phi^{+}$ of positive roots is the set $\{(1,-1),(1,2),(2,1)\}$.
@@ 2.4. Irreducible Representations
The $k$-th symmetric power $\mathop{\rm Sym}\nolimits^{k}V$ of the standard representation $V$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$ corresponds to $\lambda=(k,0)$. We denote by $V_{k}$ its lift [^3] See Rmk. 2.6 for an explanation of this choice of lift. to the irreducible representation $V_{(k,0,k,k)}$ of $\mathrm{G}_{F}$. We denote by $V^{\vee}_{k}$ its dual.
@@
We will also need to parametrize representations of $\mathrm{T}^{\prime}$. Consider the isomorphism $\mathrm{T}^{\prime}\simeq\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$ given by (10) and the natural embedding $\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}$. Fix a character $\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$, restricting under the previous embedding to the map given on ${\mathbb{Q}}$-points by
$\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)\mapsto a^{\mu}\alpha(a)^{\nu}\,,$
Then we denote by
@@
the one-dimensional $F$-vector space on which $F^{\times}$ acts via multiplication by $\chi$.
@@ 2.5. Standard $\mathbb{Q}$-Parabolic Subgroups
The only ${\mathbb{Q}}$-conjugacy class of parabolic subgroups of $\mathrm{G}$ is the class of the Borel subgroup $\mathrm{B}$. Indeed, on one hand, it is clear that $\mathrm{B}$ is a ${\mathbb{Q}}$-parabolic. On the other hand, any other parabolic has to stabilize a subspace $W$ inside $V$, hence its orthogonal $W^{\perp}$, and by dimension considerations we see that $W\cap W^{\perp}$ has to be a line. Hence any ${\mathbb{Q}}$-parabolic $\mathrm{P}$ is the stabilizer of an isotropic line, and since there is a transitive action of $\mathrm{G}({\mathbb{Q}})$ on the set of isotropic lines, $\mathrm{P}$ has to be conjugated to $\mathrm{B}$.
@@
$\mathrm{B}\simeq\mathrm{T}\rtimes\mathrm{U}$
@@ 2.6. Kostant Representatives
[34, Thm. 3.2.3] Let $\mathcal{V}_{\lambda}$ be the irreducible $\mathcal{G}$-representation of highest weight $\lambda$, and let $\delta$ be the half-sum of the positive roots of $\mathcal{G}$. Then, as $(\mathcal{P}/\mathcal{U})$-representations,
$H^{q}(\mathcal{U},\mathcal{V}_{\lambda})\simeq\bigoplus\limits_{w\in W^{\prime}|l(w)=q}\mathcal{V}^{\mathcal{P}/\mathcal{U}}_{w.(\lambda+\delta)-\delta},$
@@
(1 $(1\ 2)\star\lambda$ $=(k_{2}-1,k_{1}+1,c,r)$
(2 $(2\ 3)\star\lambda$ $=(k_{1}-k_{2}-1,-k_{2}-2,c-k_{2}-1,r)$
(1 2 $(1\ 2\ 3)\star\lambda$ $=(-k_{2}-3,k_{1}-k_{2},c-k_{2}-1,r)$
(1 3 $(1\ 3\ 2)\star\lambda$ $=(k_{2}-k_{1}-3,-k_{1}-3,c-k_{1}-2,r)$
(1 $(1\ 3)\star\lambda$ $=(-k_{1}-4,k_{2}-k_{1}-2,c-k_{1}-2,r)$
The only element of length 0 is the identity, the elements (1 2)12(1\ 2)( 1 2 ) and (2 3)23(2\ 3)( 2 3 ) have length 1, the elements (1 2 3)123(1\ 2\ 3)( 1 2 3 ) and (1 3 2)132(1\ 3\ 2)( 1 3 2 ) have length 2, and the element (1 3)13(1\ 3)( 1 3 ) has length 3.
@@
(17) $\theta:=\left(\begin{array}[]{ccc}&&-1\\ &-1&\\ -1&&\end{array}\right)$
@@ 2.7. The associated symmetric spaces and Shimura varieties
In order to describe the locally symmetric spaces associated with $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$, observe that $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Fix a standard basis of $V_{{\mathbb{R}}}$, consider the resulting embedding $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\hookrightarrow\mathop{\rm GL}\nolimits_{3}({\mathbb{C}})$ and write an element $g\in\mathrm{G}({\mathbb{R}})$ in block-matrix form as
@@
$\mathrm{S}:=\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})/K_{\infty}$
@@
$\mathrm{S}_{\Gamma}:=\Gamma\backslash\mathrm{S}$
@@
(19) $x+iy\mapsto\left(\begin{array}[]{ccc}x&0&iy\\ 0&x+iy&0\\ iy&0&x\end{array}\right)$
It defines a pure Shimura datum[24, 3.2] [^4] To make the comparison with Lan’s setting, one has to notice that $i\cdot\big{(}\begin{smallmatrix}&&1\\ &1&\\ i&&\end{smallmatrix}\big{)}\cdot J_{\beta}\cdot\big{(}\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big{)}=\big{(}\begin{smallmatrix}&&1\\ &i\beta&\\ -1&&\end{smallmatrix}\big{)}$, where the latter matrix represents a skew-hermitian form, and that conjugating the above Shimura datum by $\big{(}\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big{)}$ gives the complex conjugate of Lan’s Shimura datum (loc. cit., Eq. (3.2.5.4)).). Then, the $\mathrm{G}({\mathbb{R}})$-conjugacy class $X$ of $h$ is in canonical bijection with a disjoint union of copies of $\mathrm{S}$ and therefore acquires a canonical topology. For any compact open subgroup $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is neat in the sense of [27, Sec. 0.6], the set of double classes
@@
$\mathrm{S}_{K}({\mathbb{C}})=\mathrm{G}({\mathbb{Q}})\backslash(X\times\mathrm{G}({\mathbb{A}}_{f})/K)$
@@
Each surface $\mathrm{S}_{K}$ can be described as a moduli space of abelian threefolds equipped with additional structures depending on $K$, amongst which complex multiplication by the ring of integers of $F$. In particular, there exists an universal abelian threefold over $\mathrm{S}_{K}$.
The locally symmetric spaces $\Gamma_{i}\backslash\mathrm{S}$ are identified to the analytifications of smooth, geometrically connected quasi-projective surfaces, defined over an abelian extension of $F$ depending on $K$.

### v4

@@ 1.4. Structure of the article
In Section 2 we gather all the preliminary facts that we need about the structure of the group $\mathrm{G}$, its representation theory, its parabolic subgroups, and about its associated locally symmetric spaces, i.e., the Picard surfaces $\mathrm{S}_{K}$, together with the variations of Hodge structures carried by them.
In Section 3, we set up our notations for algebraic Hecke characters. Moreover, we recall the basic facts about their $L$-functions and their associated motives that we need to formulate the Bloch-Beilinson conjectures and to explain how Conjecture 1 follows from them.
In Section 4, we recall the structure of the Baily-Borel and Borel-Serre compactifications of Picard modular surfaces. Moreover, we revisit in detail the computation of boundary cohomology of Picard surfaces and its Hodge structure ([2]), adding the complements we need. We proceed then by recalling in detail the contents of Harder’s paper [23] on which we rely crucially: in Subsection 4.4 we record the description of boundary cohomology, using $\bar{\mathrm{S}}_{K}$, in terms of induced representations from Hecke characters, and in Section 5 we explain the description of Eisenstein cohomology, where the $L$-functions of the aforementioned Hecke characters play an essential role. We conclude the section by constructing the first piece of our extensions, the space $\partial H^{2}(\phi)$.
In Section 6, we explain how to use Rogawski’s work to construct the second piece needed for our extensions, i.e., the space $H^{2}(\phi)_{!}$. The sought-for extensions are then constructed in Section 7. The final Section 8 is devoted to showing how our extensions fit into suitable biextensions. These, as explained before, will provide the starting ground for our future work on the proof of Conjecture 1.
@@ 2. Preliminaries
The first part of this section fixes notation about the ground field $F$ and the algebraic group $\mathrm{G}$ that will be in force throughout the whole paper. In the following parts, we discuss the root system and Weyl group of $\mathrm{G}$, we fix a parametrization of its irreducible representations, and we explain the structure of its parabolic subgroups. Finally, we introduce the main geometric objects of interest: the locally symmetric spaces attached to $\mathrm{G}$, their structure of Shimura varieties, and the variations of Hodge structure associated to representations of $\mathrm{G}$.
@@ 2.1. Basic notations and conventions
The symbols ${\mathbb{A}}_{{\mathbb{K}}}$ and ${\mathbb{I}}_{{\mathbb{K}}}$ will denote respectively adèles and idèles over a number field $\mathbb{K}$} (no subscript if $\mathbb{K}$$\mathbb{Q}$}). The quadratic character of ${\mathbb{I}}_{F}$ associated by class field theory to the extension $F|{\mathbb{Q}}$ will be denoted by $\omega_{F|{\mathbb{Q}}}$.
@@
There always exists an $F$-basis of $V$ with respect to which $J$ is represented by a matrix of the form
@@
$\left(\begin{array}[]{ccc}1&&\\ &1&\\ &&-1\end{array}\right)\,.$
@@
(4) $\gamma:=\left(\begin{array}[]{ccc}\frac{1}{\sqrt{2}}&&\frac{1}{\sqrt{2}}\\ &\sqrt{\beta}&\\ \frac{1}{\sqrt{2}}&&-\frac{1}{\sqrt{2}}\end{array}\right)\,.$
@@
$\mathrm{G}(R)=\{g\in\mathop{\rm GL}\nolimits_{F\otimes_{{\mathbb{Q}}}R}(V\otimes_{{\mathbb{Q}}}R)|\ \exists\ \nu(g)\in R^{\times},J(g\cdot,g\cdot)=\nu(g)J(\cdot,\cdot)\}\,.$
@@
$\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}:=\mathop{\rm SU}\nolimits(V,J)\,.$
@@ 2.2. Structure Theory of $\mathrm{G}$
To describe the structure of $\mathrm{G}$, recall that $V_{F}$ decomposes as
@@
$V^{+}:=\{v\in V\otimes_{{\mathbb{Q}}}F\ |\ (x\otimes 1)v=(1\otimes x)v\ \forall x\in F\}\,$,
@@
(9) $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}\simeq\mathop{\rm SL}\nolimits(V_{F})\,.$
@@
and call $\mathcal{B}$} and $\mathcal{T}$ the standard (upper triangular) Borel and the maximal torus of $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathop{\rm GL}\nolimits_{3,F}$. Then,
@@
is a maximal torus that we call standard. It is isogenous to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}\times\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$. Under the isomorphism
@@
(10) $\mathrm{T}_{F}(F)\simeq\left\{(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f)|a,b,q,f\in F^{\times}\right\}\,.$
@@
$\mathrm{T}^{\prime}:=\mathrm{T}\cap\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}\,.$
It is isomorphic to $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$, and we fix the isomorphism to be the one described on $\mathbb{Q}$}-points by [^2] We adopt this convention to be coherent with [23, p. 573].
(11) $\mathrm{T}^{\prime}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)|a\in F^{\times}\right\}\,.$
@@
(12) $\mathrm{T}^{s}({\mathbb{Q}})\simeq\left\{\left(\begin{array}[]{ccc}t&&\\ &1&\\ &&t^{-1}\end{array}\right)\ |\ t\in{\mathbb{Q}}^{\times}\right\}\,.$
@@ 2.3. Root System
$k_{1}\geq k_{2}\geq 0,\ c=k_{1}+k_{2}\pmod{2},\ r=\frac{c+k_{1}+k_{2}}{2}\pmod{2}\,$.
@@
(13) $\left(\left(\begin{array}[]{ccc}a&&\\ &a^{-1}b&\\ &&b^{-1}q\end{array}\right),f\right)\mapsto a^{k_{1}-k_{2}}b^{k_{2}}q^{\frac{c-(k_{1}+k_{2})}{2}}\cdot f^{-\frac{1}{2}(r+\frac{3c-(k_{1}+k_{2})}{2})}\,.$
@@
and the group of characters of the maximal torus $\mathrm{T}_{F}^{\prime}$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$. The isomorphism is provided by the natural factorization of characters of the above shape through $\mathrm{T}^{\prime}_{F}$. We still call the resulting character $\lambda$ and denote it by $(k_{1},k_{2})$.
The root system $\Phi$ of the split semi-simple group $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}_{F}$ is of type $A_{2}$. Under the previous identifications, the system $\Delta$ of simple roots is the set $\{(1,-1),(1,2)\}$ and the set $\Phi^{+}$ of positive roots is the set $\{(1,-1),(1,2),(2,1)\}$.
@@ 2.4. Irreducible Representations
The $k$-th symmetric power $\mathop{\rm Sym}\nolimits^{k}V$ of the standard representation $V$ of $\mathrm{G}_{F}^{\mathop{\rm\tiny{der}}\nolimits}$ corresponds to $\lambda=(k,0)$. We denote by $V_{k}$ its lift [^3] See Remark 2.6 for an explanation of this choice of lift. to the irreducible representation $V_{(k,0,k,k)}$ of $\mathrm{G}_{F}$. We denote by $V^{\vee}_{k}$ its dual.
@@
We will also need to parametrize representations of $\mathrm{T}^{\prime}$. Consider the isomorphism
$\mathrm{T}^{\prime}\simeq\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$
given by (11), and the natural embedding
$\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}.$
Fix a character
$\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$
whose restriction under the above embedding is the map given on $\mathbb{Q}$}-points by
$\left(\begin{array}[]{ccc}\alpha(a)&&\\ &\alpha(a)^{-1}a&\\ &&a^{-1}\end{array}\right)\mapsto a^{\mu}\alpha(a)^{\nu}\,.$
We then denote by
@@
the one-dimensional $F$-vector space on which $F^{\times}$ acts by multiplication via $\chi$.
@@ 2.5. Standard $\mathbb{Q}$-Parabolic Subgroups
The only $\mathbb{Q}$}-conjugacy class of parabolic subgroups of $\mathrm{G}$ is the class of the Borel subgroup $\mathrm{B}$. Indeed, on one hand, it is clear that $\mathrm{B}$ is a $\mathbb{Q}$}-parabolic. On the other hand, any other parabolic has to stabilize a subspace $W$ inside $V$, hence its orthogonal $W^{\perp}$, and by dimension considerations we see that $W\cap W^{\perp}$ has to be a line. Hence, any $\mathbb{Q}$}-parabolic $\mathrm{P}$ is the stabilizer of an isotropic line, and since there is a transitive action of $\mathrm{G}({\mathbb{Q}})$ on the set of isotropic lines, $\mathrm{P}$ has to be conjugated to $\mathrm{B}$.
@@
$\mathrm{B}\simeq\mathrm{T}\rtimes\mathrm{U}\,.$
@@ 2.6. Kostant Representatives
[41, Thm. 3.2.3] Let $\mathcal{V}_{\lambda}$ be the irreducible $\mathcal{G}$-representation of highest weight $\lambda$, and let $\delta$ be the half-sum of the positive roots of $\mathcal{G}$. Then, as $(\mathcal{P}/\mathcal{U})$-representations,
$H^{q}(\mathcal{U},\mathcal{V}_{\lambda})\simeq\bigoplus\limits_{w\in W^{\prime}|l(w)=q}\mathcal{V}^{\mathcal{P}/\mathcal{U}}_{w.(\lambda+\delta)-\delta}$
@@
$(1\ 2)\star\lambda$ $=(k_{2}-1,k_{1}+1,c,r)$
$(2\ 3)\star\lambda$ $=(k_{1}-k_{2}-1,-k_{2}-2,c-k_{2}-1,r)$
$(1\ 2\ 3)\star\lambda$ $=(-k_{2}-3,k_{1}-k_{2},c-k_{2}-1,r)$
$(1\ 3\ 2)\star\lambda$ $=(k_{2}-k_{1}-3,-k_{1}-3,c-k_{1}-2,r)$
$(1\ 3)\star\lambda$ $=(-k_{1}-4,k_{2}-k_{1}-2,c-k_{1}-2,r)$
The only element of length 0 is the identity, the elements (1 2)(1\ 2) and (2 3)(2\ 3) have length 1, the elements (1 2 3)(1\ 2\ 3) and (1 3 2)(1\ 3\ 2) have length 2, and the element (1 3)(1\ 3) has length 3.
@@
(18) $\theta:=\left(\begin{array}[]{ccc}&&-1\\ &-1&\\ -1&&\end{array}\right)\,.$
@@ 2.7. The associated symmetric spaces and Shimura varieties
To describe the locally symmetric spaces associated with $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$, observe that $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\simeq\mathop{\rm SU}\nolimits(2,1)$. Fix a standard basis of $V_{{\mathbb{R}}}$, consider the resulting embedding $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})\hookrightarrow\mathop{\rm GL}\nolimits_{3}({\mathbb{C}})$ and write an element $g\in\mathrm{G}({\mathbb{R}})$ in block-matrix form as
@@
$\mathrm{S}:=\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})/K_{\infty}\,.$
@@
$\S\mathrm{S}_{\Gamma}:=\Gamma\backslash\mathrm${S}
@@
(20) $x+iy\mapsto\left(\begin{array}[]{ccc}x&0&iy\\ 0&x+iy&0\\ iy&0&x\end{array}\right)\,.$
It defines a pure Shimura datum ( [29, Sect. 3.2] [^4] To make the comparison with Lan’s setting, one has to notice that $i\cdot\big(\begin{smallmatrix}&&1\\ &1&\\ i&&\end{smallmatrix}\big)\cdot J_{\beta}\cdot\big(\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big)=\big(\begin{smallmatrix}&&1\\ &i\beta&\\ -1&&\end{smallmatrix}\big)$, where the latter matrix represents a skew-hermitian form, and that conjugating the above Shimura datum by $\big(\begin{smallmatrix}&&-i\\ &1&\\ 1&&\end{smallmatrix}\big)$ gives the complex conjugate of Lan’s Shimura datum (loc. cit., Eq. (3.2.5.4)).). Then, the $\mathrm{G}({\mathbb{R}})$-conjugacy class $X$ of $h$ is in canonical bijection with a disjoint union of copies of $\mathrm{S}$ and therefore acquires a canonical topology. For any compact open subgroup $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is neat in the sense of [34, Sect. 0.6], the set of double classes
@@
$\mathrm{S}_{K}({\mathbb{C}})=\mathrm{G}({\mathbb{Q}})\backslash(X\times\mathrm{G}({\mathbb{A}}_{f})/K)\,.$
@@
Each surface $\mathrm{S}_{K}$ can be described as a moduli space of abelian threefolds equipped with additional structures depending on $K$, amongst which complex multiplication by the ring of integers of $F$. In particular, there exists a universal abelian threefold $A_{K}$ over $\mathrm{S}_{K}$.
The locally symmetric spaces $\S\Gamma_{i}\backslash\mathrm${S} are identified with the analytifications of smooth, geometrically connected quasi-projective surfaces, defined over an abelian extension of $F$ depending on $K$.